Freddie the frog is climbing up a well. Every day he climbs up 3m but some nights he falls asleep and slips back 4m. At the start of the sixteenth day, he has climbed a total of 29m. On how many nights was he asleep?
Answer by richwmiller(17219)

15*3=45
45-29=16
16/4=4 nights he fell back